Circular No. 6248

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1995 O1 (HALE-BOPP) Z. Sekanina, Jet Propulsion Laboratory, reports: “The evolution of the

bright jet on images published on the World Wide Web since late September suggests that the general scenario that I proposed on IAUC 6223 and 6240 remains viable, with two additional refinements. The three spiral jets in late August, late September, and mid-October were almost certainly made up of dust ejecta from the same source, located in the comet’s equatorial plane (defined relative to the apparent pole of rotation at the time of emission; cf. below). It appears that, shortly after the outburst’s onset near local noon, the production of dust from this source peaks sharply, then gradually subsides, and eventually terminates at local sunset. The production peak explains the presence of the rectilinear jet in the fourth quadrant (as well as a brightness excess in the spiral feature’s westernmost portion), eliminating the need for a second source (cf. IAUC 6240). The systematic variations in the straight jet’s position angle from ∼290^{o} in the August emission episode, to 320^{o}–330^{o} in the September event, and to ∼350^{o} in the October episode (as well as similar variations in the p.a. of the easternmost tip of the spiral feature) suggest that the comet is not in the state of pure spin, unless one considers the unlikely case of an active area’s migration over the nucleus. To match the feature’s evolution during the three outbursts requires that the apparent rotation pole have moved by ∼35^{o} between late August and late September, from 40^{o} to 70^{o} in α and from 0^{o} to −20^{o} in δ (equivalent to a change from 150^{o} to 120^{o} in the apparent obliquity), and by an additional ∼20^{o} or more between late September and mid-October. The temporal separations of ∼5 weeks between the first two emission episodes, and one half this duration between the last two events, are signatures of the comet’s complex rotation state. This evidence implies a recurrence period of ∼18 days and suggests that the source failed to activate (in the sense to generate a major dust feature) in the last days of July and again in early September, thereby twice ‘skipping’ a cycle. Such ‘duds’ in cometary activity are not altogether uncommon (especially not at large r). A recurrence period of one half the proposed duration or shorter would imply too many non-events and is less likely. The apparent spin rate during the three outbursts is still unknown. The (unforeshortened) expansion velocities of the particles that made up the spiral feature in the last two episodes are found to have ranged from 30 to 50 m/s, being comparable with those in the August event (cf. IAUC 6240). Observations of additional dust features should further help refine the proposed emission scenario.”

1995 October 20 Daniel W. E. Green